OpenStudy (anonymous):

Need help! the system of equations is coincident. what are the missing values? 3x+y=7 9x+___ y=___

OpenStudy (dtan5457):

coincident means they are equal. the 2nd one has a 9x therefore the ratio is 1:3 9x+3y=21<<<missing equation

OpenStudy (bradely):

Ratio of coefficients are equal Ratio of x - coefficients = 9/3 =3 Ratio of y - coefficients = y-coefficient of second equation/1 =3 y-coefficient of second equation=1*3=3 Ratio of constants = constant of second equation/7 =3 constant of second equation =3*7 = 21 The second is 9x + 3y =21 Source: http://www.mathskey.com/question2answer/